# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v2

@@ 1 Introduction
where $\tilde{p}$ is the probability that a random NN (sampled from the ‘uniform’ prior) is equivalent to the teacher function. [^2] The proof idea is simple: the number of hypotheses sampled until a successful interpolation is $\left|\mathcal{H}\right|\lessapprox 1/\tilde{p}$. Plugging this into the standard sample complexity of a finite hypotheses class $O\left(\log\left|\mathcal{H}\right|\right)$, we obtain the result. The actual proof is slightly more complicated since $\left|\mathcal{H}\right|$ here weakly depends on the training set. Thus, to obtain generalization guarantees for NNs, we proceed to upper bound $\left(-\log\tilde{p}\right)$.
@@
In Section 7 we discuss our assumptions (teacher narrowness and weight quantization), how our results can be straightforwardly extended beyond interpolators (to functions with a non-zero training error), whether posterior sampling biases us towards sparse representations, the effect of parameterization via the minimum description length framework, and the relation of our results to SGD.
@@ 3 Generalization Bounds for Random Interpolating Hypotheses
As we show in the next result, $\tilde{p}$ plays an important role in G&C generalization. Specifically, in Appendix B.2 we derive the following generalization bound.
@@
Proof idea of Lemma 3.2 and relationship to PAC-Bayes. In Appendix B.2, we prove Lemma 3.2 by noting that the expected number of hypotheses we will consider is $1/\tilde{p}$, and so we are essentially selecting an interpolating hypothesis from the effective hypothesis class =$\mathcal{H}=\{h_{1},\ldots,h_{\tau}\}$ with $\tau\approx 1/\tilde{p}$, where the hypotheses in this class are chosen before seeing the training set. The sample complexity is thus $\log\left|\mathcal{H}\right|=\log\tau\approx-\log\tilde{p}$. The only complication is that the stopping time $\tau$ of G&C is random and depends on $\mathcal{S}$. But it is enough to bound $\tau$ very crudely (which we do with high probability), as a multiplicative factor to $\tau$ results only in an additive logarithmic factor.
@@
The same method can be straightforwardly used to extend these results to the case where the teacher NN is not a perfect interpolator, and the G&C algorithm is modified to stop when the training error is bellow some threshold instead of 0 (see Appendix B.4).
The analysis here is similar to PAC-Bayes analysis (McAllester, 1999), which also studies the behavior of a posterior over hypotheses, except that in typical PAC-Bayes analysis the bound is over the expected population error $\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]$ for a sample from the posterior. In fact, noting that $\mathrm{KL}(\mathcal{P}_{\mathcal{S}}\|\mathcal{P})=-\log\mathbb{P}_{h\sim\mathcal{P}}(\mathcal{L}_{\mathcal{S}}(h)\!=\!0)\geq-\log\tilde{p}$, a standard PAC-Bayes bound (Langford & Seeger, 2002; McAllester, 2003) will yield that with the same sample complexity as in Lemma 3.2,
@@
The difference is that Lemma 3.2 holds with high probability for a single posterior sample, instead of just in the expectation over the posterior. That is, Lemma 3.2 establishes that not only are random interpolators good on average, but only a small fraction of them are bad.
Using Markov’s inequality one can derive from (5) a high probability bound for a for a single draw from the posterior (as in Lemma 3.2), but that bound is less tight then Lemma 3.2 (see Appendix B.5). Also, a variant of the PAC-Bayes theorem that holds with high probability for a single draw from the posterior has also been presented by Alquier (2023, Theorem 2.7). As is, Alquier’s result yields a much looser bound, since it is more generic (it applies to any posterior, not just conditioning on interpolation). Thus, Lemma 3.2 can be viewed as a tighter specialization to interpolators.
A sample complexity of $(-\log\tilde{p})$ should not be surprising, and can also be obtained by an Occam Razor / Minimum Description Length learning rule $\mathrm{MDL}_{\mathcal{P}}(\mathcal{S})=\arg\max_{\mathcal{L}_{\mathcal{S}}(h)=0}\mathcal{P}(h)$ (Blumer et al. (1987), and see also Section 7.3 in Shalev-Shwartz & Ben-David (2014)). Here, we discussed how the same sample complexity is obtained by a single draw from the posterior, as in G&C. More interesting is how, starting from a uniform prior $\mathcal{P}_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ over parameters, we end up with an informative induced prior $\mathcal{P}(h)$ over hypotheses, which has high $\tilde{p}$ and thus low sample complexity. The key here is redundancy in the parameterization. In Appendix A we review the general principle of how non-uniform redundancy in the parameterization can induce non-uniform informative priors $\mathcal{P}(h)$ and thus low sample complexity. In the next sections we see how this plays specifically for NNs, analyzing $\tilde{p}$ under the prior induced by a uniform choice of NN parameters.
@@ 4 Quantized Nets Sample Complexity
Recall that the generalization bound in Lemma 3.2 depends on the effective sample complexity $\tilde{C}\triangleq-\log\left(\tilde{p}\right)$. In this section, we derive an upper bound on $\tilde{C}$ for quantized multi-layer Fully Connected (FC) Neural Networks (NNs) with a single binary output, with and without additional per-node scaling.
@@ 4.3 Extension to convolutional neural networks
Table 1 shows that our bound is consistent with a narrow teacher of non-trivial size and widths. For example, in ResNet18 the resulting channel numbers in the teacher layers are [3,8,8,16,16,32.32,64,.64]388161632.3264.64[3,8…,8,16,...16,32….32,64,....64][ 3 , 8 … , 8 , 16 , … 16 , 32 … .32 , 64 , … .64 ], where ‘3’ counts the input channels and the rest are the following hidden layers. This architecture can implement highly complex non-linear functions. Note that we cannot directly validate the existence of this teacher, as standard optimization methods may not be capable of finding such a solution efficiently, even if it exists (without over-parameterization, it is much harder to find global minima).
@@ 5 Continuous Nets Sample Complexity
Although Assumption 5.4 may be natural in some cases [^6] Realistic data many times has some intrinsic margin. For example, there is a low probability for (semantic) ‘mixing’ of dogs and birds from a distribution of natural images of dogs and birds (with a single animal in every image). , in other cases the margins $\alpha,\beta$ may decay with $N$ or with the network’s dimensions. Therefore, even if the assumption holds, for the generalization bound in Theorem 5.6 to remain meaningful, we need the margin decay rate to be sufficiently slow. [^7] For example, for Gaussian data with a random teacher it is possible to derive a lower bound in high probability on the first layer angular margin $\alpha$ and show that it decays as $1/(Nd_{0}d_{1}^{\star})$, so it adds only a multiplicative $\log(Nd_{0}d_{1}^{\star})$ factor to the bound.
@@ 7 Discussion
Beyond interpolators. In this work we focus for simplicity on NN interpolators, but in many scenarios, the training loss does not reach zero. In Appendix B.4 we show how to generalize our results to this non-realizable case, where the teacher NN does not reach zero training error (i.e., there is some irreducible error). Unfortunately, the use of the non-realizable generalization bound for finite hypothesis classes introduce a quadratic dependence of the sample complexity on the generalization error bound $\varepsilon$. It is interesting to see if this quadratic dependence can be improved.
@@ B.1 Finite Hypothesis Class PAC Generalization Bound
Here we present a classic generalization bound for finite hypothesis class adapted from (Shalev-Shwartz & Ben-David, 2014), which we will use throughout our paper.
@@
Then, for any realizable data distribution $\mathcal{D}$, with probability of at least $1-\delta$ over the choice of an i.i.d sample $S$ of size $N$, we have that for every interpolating121212(Shalev-Shwartz & Ben-David, 2014, Corollary 2.3) is more general, but we focus on interpolating hypotheses. hypothesis, $h_{\boldsymbol{\theta}}$, it holds that
@@ B.2 Proving the Generalization of Algorithm 1
[[[[Def. 3.1 extended]]]] For any hypothesis class $\mathcal{H}$, we say that $h\in\mathcal{H}$ is a teacher-equivalent w.r.t $\mathcal{D}$ (TE) model, and denote $h\equiv h^{\star}$, if
@@
If $h^{\star}\in\mathcal{H}_{\tau}$ then $\mathcal{H}_{\tau}$ is realizable, so we get from Theorem B.1 with
$N\geq\frac{\log\left(\frac{{\left\lvert{\mathcal{H}_{\tau}}\right\rvert}}{\delta_{\mathcal{S}}}\right)}{\varepsilon}=\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\varepsilon}$
@@
We now wish to upper bound the sample complexity.
@@
$\log\left(\frac{\log\left(\delta_{h}\right)+\log\left(1-\tilde{p}\right)}{\delta_{\mathcal{S}}\log\left(1-\tilde{p}\right)}\right)$ $\leq\log\left(\frac{\log\left(\delta_{h}\right)-\tilde{p}}{\delta_{\mathcal{S}}\left(-\tilde{p}-{p^{\star}}^{2}\right)}\right),$
@@
$\log\left(\frac{\log\left(\delta_{h}\right)+\log\left(1-\tilde{p}\right)}{\delta_{\mathcal{S}}\log\left(1-\tilde{p}\right)}\right)\leq\log\left(\frac{1}{\delta_{\mathcal{S}}\left(\tilde{p}+{p^{\star}}^{2}\right)}\right)+\log\left(\tilde{p}+\log\left(\frac{1}{\delta_{h}}\right)\right)\,.$ (17)
We now recall that $\log(c+x)$ is a concave function, and therefore its graph in any $x$ is below the graph of its tangent in $x=0$, that is, from (15),
@@
$\log\left(\frac{1}{\tilde{p}+{p^{\star}}^{2}}\right)\leq\log\left(\frac{1}{\tilde{p}}\right)\,,$ (20)
and therefore combining (20), (19) and (17) together with 14 we finish the proof:
@@
$\leq\log\left(\frac{1}{\delta_{\mathcal{S}}\left(\tilde{p}+{p^{\star}}^{2}\right)}\right)+\log\left(\tilde{p}+\log\left(\frac{1}{\delta_{h}}\right)\right)$
@@
We are now ready to prove Theorem B.2:
Proof.
(of Theorem B.2) Note that $h\sim\mathcal{P}_{\mathcal{S}}$ is interpolating by definition, and therefore, from Lemma B.6, we have that for any $\tau\in\mathbb{N}$ it holds that
@@
Where $\bar{h}\left(S\right)\in\mathcal{H}_{\tau}$ is some arbitrary interpolate, since from Lemma B.6 this probability is equal for any interpolator. For $\tau=\left\lceil\frac{\log\left(\delta_{h}\right)}{\log\left(1-\tilde{p}\right)}\right\rceil$ we similarly get
@@
$=1-\delta_{\mathcal{S}}+\frac{\delta^{2}}{4}$
@@
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\right]$
and then we use Markov’s inequality
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\geq\delta\right)\leq\frac{\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\right]}{\delta}\leq\delta\,,$

### v3

@@ 1 Introduction
where $\tilde{p}$ is the probability that a random NN (sampled from the ‘uniform’ prior) is equivalent to the teacher function. [^2] This is a generic argument and essentially a special case of a PAC-Bayes guarantee, but we give a simple proof based on the following idea: the number of hypotheses sampled until a successful interpolation is $\left|\mathcal{H}\right|\lessapprox 1/\tilde{p}$. Plugging this into the standard sample complexity of a finite hypotheses class $O\left(\log\left|\mathcal{H}\right|\right)$, we obtain the result. The actual proof is slightly more complicated since $\left|\mathcal{H}\right|$ here weakly depends on the training set. Thus, to obtain generalization guarantees for NNs, we proceed to upper bound $\left(-\log\tilde{p}\right)$.
@@
In Section 7 we discuss our assumptions (teacher narrowness and weight quantization), how our results can be straightforwardly extended beyond interplators (to functions with a non-zero training error), whether posterior samping biases us towards sparse represetnations, the effect of parameterization via the minimum description length framework, and the relation of our results to SGD.
@@ 3 Generalization Bounds for Random Interpolating Hypotheses
As we show in the next result, $\tilde{p}$ plays an important role in G&C generalization. Specifically, in Appendix B.2 we prove the following generalization bound
@@
Proof idea of Lemma 3.2. In Appendix B.2, we provide a self-contained proof of Lemma 3.2 by noting that the expected number of hypotheses we will consider is $1/\tilde{p}$, and so we are essentially selecting an interpolating hypothesis from the effective hypothesis class =$\mathcal{H}=\{h_{1},\ldots,h_{\tau}\}$ with $\tau\approx 1/\tilde{p}$, where the hypotheses in this class are chosen before seeing the training set. The sample complexity is thus $\log\left|\mathcal{H}\right|=\log\tau\approx-\log\tilde{p}$. The only complication is that the stopping time $\tau$ of G&C is random and depends on $\mathcal{S}$. But it is enough to bound $\tau$ very crudely (which we do with high probability), as a multiplicative factor to $\tau$ results only in an additive logarithmic factor.
@@
A similar result can also be proved using PAC-Bayes, as we discuss later in this section. Lastly, the same method can be straightforwardly used to extend these results to the case where the teacher NN is not a perfect interpolator, and the G&C algorithm is modified to stop when the training error is below some threshold instead of 0 (see Appendix B.4).
Relationship to PAC-Bayes The analysis here may also be seen as a special case of PAC-Bayes analysis (McAllester, 1999), which studies the behavior of a “posterior” over hypotheses in terms of its KL-divergence to the prior, where here we specialize to a specific posterior $\mathcal{P}_{\mathcal{S}}$, conditioning on interpolation.
Noting that $\mathrm{KL}(\mathcal{P}_{\mathcal{S}}\|\mathcal{P})=-\log\mathbb{P}_{h\sim\mathcal{P}}(\mathcal{L}_{\mathcal{S}}(h)\!=\!0)\geq-\log\tilde{p}$, a standard PAC-Bayes bound (Langford & Seeger, 2002; McAllester, 2003) will yield that with the same sample complexity as in Lemma 3.2,
@@
The subtlety is that (5) bounds the expected population error $\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]$ for a sample $h\sim\mathcal{P}_{\mathcal{S}}$ from the posterior, while Lemma 3.2 holds with high probability for a single posterior sample. That is, Lemma 3.2 establishes that not only are random interpolators good on average, but only a small fraction of them are bad [^4] Using Markov’s inequality one can derive from (5) a high probability bound for a single draw from the posterior, but with a sample complexity that depends polynomially rather than logarithmically on the failure probability $\delta$ (see Appendix B.5)..
High-probability PAC-Bayes guarantees on a single draw from the posterior, as in Lemma 3.2, have been derived by Catoni (2007) and simplified by (Alquier, 2023). Applying Alquier (2023, Theorem 2.7), yields a single-posterior-sample guarantee as in Lemma 3.2, but with a sample complexity of $\Theta\left(\tfrac{-\log(\tilde{p})+\log(1/\delta)}{\epsilon^{2}}\right)$ that depends quadratically $1/\epsilon$. The reason is that Alquier’s analysis is more generic and applies also to posteriors with high empirical error. Lemma 3.2 can be viewed as a tighter specialization to interpolators, thus allowing a rate of $1/\epsilon$, as in McAllester’s bounds for the average-over-the-posterior. It should be possible to write down a generic PAC-Bayes guarantee that both applies to single draws from the posterior and yields a $1/\epsilon$ rate for interpolators and then derive Lemma 3.2 from it. However, for the sake of completeness, and to give a simplified intuition, we choose instead to present a self-contained specialized proof of the Lemma in Appendix B.2.
Occam and Redundancy A sample complexity of $(-\log\tilde{p})$ should not be surprising, and can also be obtained by an Occam Razor / Minimum Description Length learning rule $\mathrm{MDL}_{\mathcal{P}}(\mathcal{S})=\arg\max_{\mathcal{L}_{\mathcal{S}}(h)=0}\mathcal{P}(h)$ (Blumer et al. (1987), and see also Section 7.3 in Shalev-Shwartz & Ben-David (2014)). Here, we discussed how the same sample complexity is obtained by a single draw from the posterior, as in G&C. More interesting is how, starting from a uniform prior $\mathcal{P}_{\boldsymbol{\theta}}(\boldsymbol{\theta})$ over parameters, we end up with an informative induced prior $\mathcal{P}(h)$ over hypotheses, which has high $\tilde{p}$ and thus low sample complexity. The key here is redundancy in the parameterization. In Appendix A we review the general principle of how non-uniform redundancy in the parameterization can induce non-uniform informative priors $\mathcal{P}(h)$ and thus low sample complexity. In the next sections we see how this plays specifically for NNs, analyzing $\tilde{p}$ under the prior induced by a uniform choice of NN parameters.
@@ 4 Quantized Nets Sample Complexity
The generalization bound in Lemma 3.2 depends on the effective sample complexity $\tilde{C}\triangleq-\log\left(\tilde{p}\right)$. In this section, we derive an upper bound on $\tilde{C}$ for quantized multi-layer Fully Connected (FC) Neural Networks (NNs) with a single binary output, with and without additional per-node scaling.
@@ 4.3 Extension to convolutional neural networks
Table 1 shows that our bound is consistent with a narrow teacher of non-trivial size and widths. For example, in ResNet18 the resulting channel numbers in the teacher layers are $[3,8,...,8,16,...,16,32,...,32,64,....64]$, where ‘3’ counts the input channels and the rest are the following hidden layers. This architecture can implement highly complex non-linear functions. Note that we cannot directly validate the existence of this teacher, as standard optimization methods may not be capable of finding such a solution efficiently, even if it exists (without over-parameterization, it is much harder to find global minima).
@@ 5 Continuous Nets Sample Complexity
Although Assumption 5.4 may be natural in some cases [^7] Realistic data many times has some intrinsic margin. For example, there is a low probability for (semantic) ‘mixings’ of dogs and birds from a distribution of natural images of dogs and birds (with a single animal in every image). , in other cases the margins $\alpha,\beta$ may decay with $N$ or with the network’s dimensions. Therefore, even if the assumption holds, for the generalization bound in Theorem 5.6 to remain meaningful, we need the margin decay rate to be sufficiently slow. [^8] For example, for Gaussian data with a random teacher it is possible to derive a lower bound in high probability on the first layer angular margin $\alpha$ and show that it decays as $1/(Nd_{0}d_{1}^{\star})$, so it adds only a multiplicative $\log(Nd_{0}d_{1}^{\star})$ factor to the bound.
@@ 7 Discussion
Beyond interpolators. In this work we focus for simplicity on NN interpolators, but in many scenarios, the training loss does not reach zero. In Appendix B.4 we show how to generalize our results to this non-realizable case, where the teacher NN does not reach zero trainning error (i.e., there is some irreducible error). Unfortunately, the use of the non-realizable generalization bound for finite hypothesis classes introduce a quadratic dependence of the sample complexity on the generalization error bound $\varepsilon$. It is interesting to see if this quadratic dependence can be improved.
@@ B.1 Finite Hypothesis Class PAC Generalization Bound
Here we present a classic generalization bound for finite hypothesis classes adapted from (Shalev-Shwartz & Ben-David, 2014), which we will use throughout our paper.
@@
Then, for any realizable data distribution $\mathcal{D}$, with probability of at least $1-\delta$ over the choice of an i.i.d sample $S$ of size $N$, we have that for every interpolating131313(Shalev-Shwartz & Ben-David, 2014, Corollary 2.3) is more general, but we focus on interpolating hypotheses. hypothesis, $h_{\boldsymbol{\theta}}$, it holds that
@@ B.2 Proving the Generalization of Algorithm 1
[[[[Def. 3.1 extended]]]] For any hypothesis class $\mathcal{H}$, we say that $h\in\mathcal{H}$ is a teacher-equivalent (TE) model w.r.t $\mathcal{D}$, and denote $h\equiv h^{\star}$, if
@@
If $h^{\star}\in\mathcal{H}_{\tau}$ then $\mathcal{H}_{\tau}$ is realizable, so from Theorem B.1 with
$N\geq\frac{\log\left(\frac{{\left\lvert{\mathcal{H}_{\tau}}\right\rvert}}{\delta_{\mathcal{S}}}\right)}{\varepsilon}=\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\varepsilon}\,,$
@@
We now wish to bound the sample complexity.
@@
$\log\left(\frac{\log\left(\delta_{h}\right)+\log\left(1-\tilde{p}\right)}{\delta_{\mathcal{S}}\log\left(1-\tilde{p}\right)}\right)$ $\leq\log\left(\frac{\log\left(\delta_{h}\right)-\tilde{p}}{\delta_{\mathcal{S}}\left(-\tilde{p}-{\tilde{p}}^{2}\right)}\right),$
@@
$\log\left(\frac{\log\left(\delta_{h}\right)+\log\left(1-\tilde{p}\right)}{\delta_{\mathcal{S}}\log\left(1-\tilde{p}\right)}\right)\leq\log\left(\frac{1}{\delta_{\mathcal{S}}\left(\tilde{p}+{\tilde{p}}^{2}\right)}\right)+\log\left(\tilde{p}+\log\left(\frac{1}{\delta_{h}}\right)\right)\,.$ (17)
We now recall that $\log(c+x)$ is a concave function, and therefore its graph in any $x$ is below the graph of its tangent at $x=0$, that is, from (15),
@@
$\log\left(\frac{1}{\tilde{p}+{\tilde{p}}^{2}}\right)\leq\log\left(\frac{1}{\tilde{p}}\right)\,,$ (20)
and therefore combining (20), (19) and (17) together with (14) we finish the proof:
@@
$\leq\log\left(\frac{1}{\delta_{\mathcal{S}}\left(\tilde{p}+{\tilde{p}}^{2}\right)}\right)+\log\left(\tilde{p}+\log\left(\frac{1}{\delta_{h}}\right)\right)$
@@
We are now ready to prove Theorem B.2.
Proof of Theorem B.2.
Note that $h\sim\mathcal{P}_{\mathcal{S}}$ is interpolating by definition, and therefore, from Lemma B.6, we have that for any $\tau\in\mathbb{N}$ it holds that
@@
Where $\bar{h}\left(S\right)\in\mathcal{H}_{\tau}$ is some interpolator. For $\tau=\left\lceil\frac{\log\left(\delta_{h}\right)}{\log\left(1-\tilde{p}\right)}\right\rceil$ we similarly get
@@
$=1-\delta+\frac{\delta^{2}}{4}$
@@
$=\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\right]\,,$
so using Markov’s inequality we have
$\mathbb{P}_{\mathcal{S}}\left(\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\geq\delta\right)\leq\frac{\mathbb{E}_{\mathcal{S}\sim\mathcal{D}^{N}}\left[\mathbb{P}_{h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\geq\varepsilon\right)\right]}{\delta}\leq\delta\,.$